# Turbulence and far-from-equilibrium equation of state of Bogoliubov waves in Bose-Einstein Condensates

### v2

@@ Acknowledgements.
This work was funded by the Simons Foundation Collaboration grant Wave Turbulence (Award ID 651471). This work was granted access to the high-performance computing facilities under GENCI (Grand Equipement National de Calcul Intensif) A0102A12494 (IDRIS and ADASTRA), the OPAL infrastructure from Université Côte d’Azur, supported by the French government, through the UCAJEDI Investments in the Future project managed by the National Research Agency (ANR) under Reference No. ANR-15-IDEX-01.
Turbulence and far-from-equilibrium equation of state of Bogoliubov waves in Bose-Einstein Condensates: Supplemental material
@@ A.1 Hamiltonian formation and wave interaction amplitude $V^{1}_{23}$
[…]
@@
where $\omega_{k}=c_{\text{s}}k\sqrt{1+(k\xi)^{2}/2}$ is the Bogoliubov dispersion relation, and $\delta_{1,2,3}$ is $1$ if ${\bf k}_{1}+{\bf k}_{2}+{\bf k}_{3}=0$, and 00 otherwise.
@@ A.2 A different method to obtain the interaction coefficient: the Bogoliubov transformation
[…]
@@ A.3 Angular average of $\delta({\bf k}-{\bf k}_{1}-{\bf k}_{2})$
Note that for expression (A18) to be determined, the vectors ${\bf k},{\bf k}_{1}$ and ${\bf k}_{2}$ cannot be strictly co-linear, as is the case for non-dispersive sound. Therefore, in this derivation, it is essential that the system is dispersive. Otherwise, $\Delta$ is $/0$ 0 undetermined, i.e. the above method originated from Zakharov and Sagdeev (1970) fails, as was already stressed in Zakharov and Sagdeev (1970).
@@
For the degenerate resonant manifold of non-dispersive sound waves, we have ${\bf k}||{\bf k}_{1}||{\bf k}_{2}$, $\operatorname{sign}(k_{1}+k_{2}-k)=0$, so we get
@@ B.1 Zakharov-Sagdeev spectrum
where $f_{1}(q,x)=\left(q(1-q)\right)^{2-x}\left(1-q^{x}-(1-q)^{x}\right)$ and $f_{2}(q,x)=\left(q(1+q)\right)^{2-x}\left((1+q)^{x}-q^{x}-1\right)$. Note that $f_{1}(q,x)-f_{2}(q,x)\propto q^{4-x}$ when $q\to 0$ and $f_{2}(q,x)\propto q^{3-x}$ when $q\to\infty$, which give us a convergence interval for $I(x)$, namely $54<x<$ 5. $I(x)$ is also convergent at $x=1$ with $I(1)=0$ that corresponds to the thermal equilibrium solution.
@@ B.2 Kolmogov-Zakharov spectrum for short Bogoliubov waves
Note that $g_{1}(q,x)\,,g_{2}(q,x)\propto q^{3-x}$ when $q\to 0$, and $g_{2}(q,x)\propto q^{1-2x}$ when $q\to\infty$, we find that $I(x)$ is convergent for $41<x<$ 4. Using Mathematica, we get the analytical result:
@@ C.1 Numerical simulations of the wave kinetic equation
We integrate the WKE (A26) using the code WavKinS Krstulovic and Labarre (2024). The code uses a logarithmic mesh $\{k_{i}=C\lambda^{i},i=1,\ldots,N_{r}\}$, where $\lambda$ is chosen in order to span the integration domain $[k_{\rm min},k_{\rm max}]$. Intermesh values are obtained using a linear interpolation that ensures positivity. Integrals are performed using the trapezoidal rule, after performing a change of variable in the following manner
@@ C.2 Numerical parameters
1 Bogoliubov WKE $32768$ $1\times 10^{-2}/\xi$ $200/\xi$ $1.032\times 10^{14}$ 6 4 0 $1.5\times 10^{-2}$ 1 80
2 acoustic WKE 8196 $1\times 10^{-4}$ $1$ $2.419\times 10^{16}$ 2 2 0.01 0.028 1 0.6
@@ C.3 Verifying the assumptions of the Wave Turbulence theory for the GPE simulations
The nonlinearity of the systems is estimated by measuring the frequency broadening $\delta\omega(k)$ around $\tilde{\omega}(k)$ such that $S(\omega,k)$ gets its maximum value for each fixed $k$. We plot the frequency broadening of run 3 and run 4 in Fig. S3(c) and (d), respectively. The WWT theory requires that: 1) The linear frequency is much bigger than nonlinear frequency broadening, $\omega_{\text{Bogol}}(k)>\delta\omega(k)$; 2) A stronger requirement for the weakly-dispersive waves Zakharov et al. (1992): the dispersion correction to non-dispersive frequency, $\omega_{\text{disp}}(k)=c_{\text{s}}k(\sqrt{1+k^{2}\xi^{2}/2}-1)$, is greater than $\delta\omega(k)$; 3) To trigger the interactions, the nonlinear broadening must be greater than the spacing between the eigenmodes of the periodic box, $\delta\omega(k)>\Delta\omega_{\text{Bogol}}(k)$, where $\Delta\omega_{\text{Bogol}}(k)=\Delta k\,\mathrm{d}\omega_{\text{Bogol}}(k)/\mathrm{d}k$ and $\Delta k=2\pi/L$ is the $k$-grid spacing. According to Fig. S3(c), $10.2<k\xi<$ 1 meets the three conditions and it is also the range where the energy spectrum obtained by GPE run 3 agrees well with the one obtained by WKE run 1, as shown in Fig. 1. As noted, Kochurin and Kuznetsov (2024) obtained ZS spectrum in non-dispersive acoustic waves, which means condition 2) might not be necessary (this complicated issue is far from being settled). WWT for short waves requires that $\omega_{\text{short}}(k)>\delta\omega(k)$ and $\delta\omega(k)>\Delta\omega_{\text{short}}(k)$, where $\Delta\omega_{\text{short}}(k)=\Delta k\,\mathrm{d}\omega_{\text{short}}(k)/\mathrm{d}k$. The $k\xi$-range that satisfies the two constraints of WTT, as shown in Fig. S3(d), are 2-15 for run 4. This range is consistent with the one where expected KZ spectra are observed in the main text. Details on how the normalized spatio-temporal spectral density and frequency broadening are calculated can be found in Zhu et al. (2022).
@@ D.1 Equation of state for the 4-wave regime
In reference Zhu et al. (2023a), the constant energy flux solution in the $4$-wave regime was derived analytically and confirmed numerically. Its $\log$-corrected spectrum reads

### v4

@@ Acknowledgements.
This work was funded by the Simons Foundation Collaboration grant Wave Turbulence (Award ID 651471). This work was provided with computing HPC and storage resources of IDRIS and CINES under the allocations A0152A14637 and A0172A14637 made by GENCI on the Jean Zay (SKL partition) and Adastra (GENOA partition) supercomputers. This work was granted access to the OPAL infrastructure from Université Côte d’Azur, supported by the French government, through the UCAJEDI Investments in the Future project managed by the National Research Agency (ANR) under Reference No. ANR-15-IDEX-01.
@@ A.1 Hamiltonian formation and wave interaction amplitude $V^{1}_{23}$
$\mathcal{S}=\frac{1}{\rho_{0}L^{3}}\int dtd^{3}x\left[-\rho\dot{\phi}-\frac{\rho}{2}(\nabla\phi)^{2}-\frac{c_{s}^{2}}{2\rho_{0}}(\rho-\rho_{0})^{2}-c_{\text{s}}^{2}\xi^{2}\left(\nabla\sqrt{\rho}\right)^{2}\right],$ (A1)
@@
where $\omega_{k}=c_{\text{s}}k\sqrt{1+(k\xi)^{2}/2}$ is the Bogoliubov dispersion relation, and $\delta_{1,2,3}$ is $1$ if ${\bf k}_{1}+{\bf k}_{2}+{\bf k}_{3}=0$, and 0 otherwise.
@@ A.2 A different method to obtain the interaction coefficient: the Bogoliubov transformation
$H_{2}=\sum_{\bf k}c_{\text{s}}^{2}(1+k^{2}\xi^{2})|\psi_{\bf k}|^{2}+\frac{c_{\text{s}}^{2}}{2}(\psi_{\bf k}\psi_{-{\bf k}}+\psi^{*}_{\bf k}\psi^{*}_{-{\bf k}})\,,\quad H_{3}=\sum_{1,2,3}c_{\text{s}}^{2}\psi_{{\bf k}_{1}}\psi^{*}_{{\bf k}_{2}}\psi^{*}_{{\bf k}_{3}}\delta^{1}_{23}+\text{c.c}\,.$ (A10)
@@ A.3 Angular average of $\delta({\bf k}-{\bf k}_{1}-{\bf k}_{2})$
Note that for expression (A18) to be determined, the vectors ${\bf k},{\bf k}_{1}$ and ${\bf k}_{2}$ cannot be strictly co-linear, as is the case for non-dispersive sound. Therefore, in this derivation, it is essential that the system is dispersive. Otherwise, $\Delta$ is $0/0$ undetermined, i.e. the above method originated from Zakharov and Sagdeev (1970) fails, as was already stressed in Zakharov and Sagdeev (1970).
@@
[…]
@@ B.1 Zakharov-Sagdeev spectrum
where $f_{1}(q,x)=\left(q(1-q)\right)^{2-x}\left(1-q^{x}-(1-q)^{x}\right)$ and $f_{2}(q,x)=\left(q(1+q)\right)^{2-x}\left((1+q)^{x}-q^{x}-1\right)$. Note that $f_{1}(q,x)-f_{2}(q,x)\propto q^{4-x}$ when $q\to 0$ and $f_{2}(q,x)\propto q^{3-x}$ when $q\to\infty$, which give us a convergence interval for $I(x)$, namely $4<x<5$. $I(x)$ is also convergent at $x=1$ with $I(1)=0$ that corresponds to the thermal equilibrium solution.
@@ B.2 Kolmogov-Zakharov spectrum for short Bogoliubov waves
Note that $g_{1}(q,x)\,,g_{2}(q,x)\propto q^{3-x}$ when $q\to 0$, and $g_{2}(q,x)\propto q^{1-2x}$ when $q\to\infty$, we find that $I(x)$ is convergent for $1<x<4$. Using Mathematica, we get the analytical result:
@@ C.1 Numerical simulations of the wave kinetic equation
We integrate the WKE (A26) using the code WavKinS Krstulovic and Labarre (2024). The code uses a logarithmic mesh $\{k_{i}=C\lambda^{i},i=1,\ldots,N_{r}\}$, where $\lambda$ is chosen in order to span the integration domain $k_{\rm min},k_{\rm max}]$. Intermesh values are obtained using a linear interpolation that ensures positivity. Integrals are performed using the trapezoidal rule, after performing a change of variable in the following manner
@@ C.2 Numerical parameters
1 Bogoliubov WKE 3276832768 $1\times 10^{-2}/\xi$ $200/\xi$ $1.032\times 10^{14}$ 6 4 0 $1.5\times 10^{-2}$ 1 80
2 acoustic WKE 8196 $\times 10^{-4} 11$ $419\times 10^{16} 2 2$ 0.01 0.028 1 0.6
@@ C.3 Verifying the assumptions of the Wave Turbulence theory for the GPE simulations
The nonlinearity of the systems is estimated by measuring the frequency broadening $\delta\omega(k)$ around $\tilde{\omega}(k)$ such that $S(\omega,k)$ gets its maximum value for each fixed $k$. We plot the frequency broadening of run 3 and run 4 in Fig. S3(c) and (d), respectively. The WWT theory requires that: 1) The linear frequency is much bigger than nonlinear frequency broadening, $\omega_{\text{Bogol}}(k)>\delta\omega(k)$; 2) A stronger requirement for the weakly-dispersive waves Zakharov et al. (1992): the dispersion correction to non-dispersive frequency, $\omega_{\text{disp}}(k)=c_{\text{s}}k(\sqrt{1+k^{2}\xi^{2}/2}-1)$, is greater than $\delta\omega(k)$; 3) To trigger the interactions, the nonlinear broadening must be greater than the spacing between the eigenmodes of the periodic box, $\delta\omega(k)>\Delta\omega_{\text{Bogol}}(k)$, where $\Delta\omega_{\text{Bogol}}(k)=\Delta k\,\mathrm{d}\omega_{\text{Bogol}}(k)/\mathrm{d}k$ and $\Delta k=2\pi/L$ is the $k$-grid spacing. According to Fig. S3(c), $0.2<k\xi<1$ meets the three conditions and it is also the range where the energy spectrum obtained by GPE run 3 agrees well with the one obtained by WKE run 1, as shown in Fig. 1. As noted, Kochurin and Kuznetsov (2024) obtained ZS spectrum in non-dispersive acoustic waves, which means condition 2) might not be necessary (this complicated issue is far from being settled). WWT for short waves requires that $\omega_{\text{short}}(k)>\delta\omega(k)$ and $\delta\omega(k)>\Delta\omega_{\text{short}}(k)$, where $\Delta\omega_{\text{short}}(k)=\Delta k\,\mathrm{d}\omega_{\text{short}}(k)/\mathrm{d}k$. The $k\xi$-range that satisfies the two constraints of WTT, as shown in Fig. S3(d), are 2-15 for run 4. This range is consistent with the one where expected KZ spectra are observed in the main text. Details on how the normalized spatio-temporal spectral density and frequency broadening are calculated can be found in Zhu et al. (2022).
@@ D.1 Equation of state for the 4-wave regime
In reference Zhu et al. (2023), the constant energy flux solution in the $4$-wave regime was derived analytically and confirmed numerically. Its $\log$-corrected spectrum reads